# Balancing Passenger Transport and Power Distribution: A Distributed Dispatch Policy for Shared Autonomous Electric Vehicles

### v1

@@ III-B Convergence
The relationship between variables in a Markovian structure (such as the queuing model used in vehicle dispatch) can make the convergence of the ADMM algorithm even harder. In general, for a single vehicle, an infeasibility at the end of the time horizon could require the entire route to be altered in the next iteration. However, the benefit of the model predictive control algorithm is that solutions at timesteps in the future are discarded. Thus, it is less important to achieve either optimality or feasibility at future times (particularly if future energy and passenger demands are not well known). For this reason, it is helpful to modify the penalty parameters $\rho$ to be time-dependent. We define $\rho(t)=\frac{\rho}{(\tau+1)^{\alpha}}$, where $\alpha$ is a small constant. This reduces the ADMM penalty over time, making large shifts to vehicle routing early in the time horizon less likely.
@@ IV Numerical Demonstrations
In this paper, we examine the properties of our dispatch strategy on a small but realistic network (5-node distribution system and 5-node transportation system with 10 SAEVs found in Robbennolt [19]). This allows for comparison between the centralized solution and the decentralized ADMM solution. To solve the centralized problems and the sub-problems of the decentralized problem we use IBM ILOG CPLEX version 12.9. Additional tests are run on the larger Sioux Falls network (24 nodes) [33] using the IEEE-85 node network for the electric distribution system [34] with 150 SAEVs. A simulation model was created in Java to test to long term behavior of vehicles and collect queuing and energy service information. Simulations were run on a laptop computer with an Intel Core i7-1165 at 2.80 GHz and 16 GB of RAM.
Passengers are loaded based on a Poisson distribution and electric demands are perturbed around the mean using a normal distribution (within 10%). We consider a contingency scenario in which approximately 35% (21% for the IEEE 85-node network) of the energy demand cannot be served from the grid and must be moved by vehicles. We consider vehicles acting as SAVs that do not discharge to the grid (but still must charge from the grid), TESSs that charge and discharge but do not serve passengers, and SAEVs dispatched using policy $\pi^{\star}$ (showcased in Figures 4 and 5). For the toy network, we also consider a scenario in which both types of vehicles are present, half SAVs and half TESSs. For this case study all of the pricing functions $\mu$ are given a constant value except for $\mu_{1}$ ($\mu_{0}=1$/pass, $\mu_{2}=500$/MWh, $\mu_{3}=100$/MWh, and $\mu_{4}=50$/MWh). We modify $\mu_{1}$ to be discounted to encourage vehicles to be served more quickly: $\mu_{1}=20-\frac{0.1}{\tau+1}$/hr.
@@
We next compare our decentralized ADMM algorithm with the centralized version using CPLEX. Examining the SAEV dispatch policy over the first three hours of operation in Figures 6 and 7, we can observe the process of queue formation and energy service. After three hours, both policies have stable passenger queues and serve approximately the same portion of the energy demand over time. These plots show that the decentralized policy is able to approximate the optimal solution, without high levels of communication. The passenger queues stabilize to similar queue lengths, and the energy service is only slightly lower. Additional heuristics built into the end of the process, ensuring vehicles are discharging at higher rates to serve all demand could be explored in the future to speed the convergence of the algorithm and reduce the difference between the two curves in Figure 7.
@@
We run similar tests on the larger Sioux Falls transportation network with the IEEE 85-node distribution network. These networks are shown in Figure 8 and are connected by four sets of charging stations. We test the operations of vehicles when used as SAV, TESSs, and SAEVs and there are two breaks in the electric grid between nodes 31–32 and nodes 34–44. Centralized solutions take hours (if they can be found at all) using the centralized problem in CPLEX, while good solutions can be found in 2–5 minutes using the decentralized ADMM approach (assuming vehicles can optimize their own routes in parallel). Additional speed gains can be found in the future by developing heuristics for the lower-level routing problems.
@@
Figures 9 and 10 show a very similar pattern as for the smaller test network. The SAEVs are able to stabilize the demand with only slightly longer queues than if they were operated as SAVs, while also providing power to impacted areas. This is because vehicles already traveling between nodes carrying passengers can stop for a short time to discharge before carrying passengers the other direction and recharging. While neither fleet stabilizes passenger demands in the first 3 hours, the SAEV dispatch shows only slightly longer queues than that SAV fleet.
@@
These case studies suggest several implications for practical applications of SAEV routing methodologies and for future algorithmic improvements. This study suggests that ignoring interactions between SAEVs and the grid can be detrimental as charging and discharging behaviors add additional demands on the time of these vehicles. Ignoring such interactions could lead to underestimates of queue lengths as shown in Figures 4 and 9. On the other hand, ignoring the need to serve passengers (particularly during disaster scenarios) could lead to overestimates of the amount of power SAEVs can carry across outages (see figures 5 and 10). Though this paper focused on disaster scenarios, vehicles may also use vehicle to grid charging to reduce demand fluctuations, provide peak shaving, or many other grid services. However, demand fluctuations of energy and vehicles are linked, so this behavior needs to be better analyzed in a comprehensive framework to understand these impacts.
@@ V Conclusions
Future work should develop heuristics or use learning-based optimizers to solve the problem more quickly (for real-time implementation on realistic networks). The lower-level problems of the ADMM heuristic were solved with CPLEX, though many exhibit structures that could be exploited in additional exact or heuristic solution approaches. Additional research should also examine the sensitivity to the cost functions, fleet size, charging/discharging rate, length of the time horizon, and behavior under different loading scenarios (i.e. more realistic demand profiles). Similarly, additional objective functions should be tested for day-to-day operation that could minimize demand fluctuations, prioritize green energy sources, incorporate equity concerns, etc. In addition to better understanding the SAEV routing problem, we believe that our proposed ADMM heuristic could be applicable to other vehicle routing approaches. This subproblem was shown to be very effective when the network was congested, and variants could be applied to other large-scale dispatch and routing problems.

### v2

@@ III-B Convergence
The relationship between variables in a Markovian structure (such as the queuing model used in vehicle dispatch) can make the convergence of the ADMM algorithm harder. In general, for a single vehicle, an infeasibility at the end of the time horizon could require the entire route to be altered in the next iteration. However, in the model predictive control algorithm, solutions at future timesteps are discarded. Thus, it is less important to achieve either optimality or feasibility at future times (particularly if future energy and passenger demands are not well known). For this reason, it is helpful to modify the penalty parameters $\rho$ to be time-dependent. We define $\rho(t)=\frac{\rho}{(\tau+1)^{\alpha}}$, where $\alpha$ is a small constant. This reduces the ADMM penalty over time, making large shifts to vehicle routing early in the time horizon less likely.
@@ IV Numerical Demonstrations
In this paper, we examine the properties of our dispatch strategy on a small but realistic network (5-node distribution system and 5-node transportation system with 10 SAEVs found in Robbennolt [34]). This allows for comparison between the centralized solution and the decentralized ADMM solution. To solve the centralized problem (MILP) and the sub-problems of the decentralized problem (all MILPs or MIQPs) we use IBM ILOG CPLEX version 12.9. Additional tests are run on the larger Sioux Falls network (24 nodes) [44] using the IEEE-85 node network for the electric distribution system [45] with 150 SAEVs. A simulation model was created in Java to test to long term behavior of vehicles and collect queuing and energy service information. Simulations were run on a laptop computer with an Intel Core i7-1165 at 2.80 GHz and 16 GB of RAM.
Passengers are loaded based on a Poisson distribution and electric demands are perturbed around the mean using a normal distribution (within 10%). We consider a contingency scenario in which approximately 35% (21% for the IEEE 85-node network) of the energy demand cannot be served from the grid and must be moved by vehicles. We consider vehicles acting as SAVs that do not discharge to the grid (but must charge from the grid), TESSs that charge and discharge but do not serve passengers, and SAEVs dispatched using policy $\pi^{\star}$ (showcased in Figures 4 and 5). For the toy network, we also consider a scenario in which both types of vehicles are present, half SAVs and half TESSs. For this case study all of the pricing functions $\mu$ are given a constant value ($\mu_{0}=1$/pass, $\mu_{1}=20$/hr, $\mu_{2}=500$/MWh, $\mu_{3}=100$/MWh, and $\mu_{4}=50$/MWh).
@@
We next compare our decentralized ADMM algorithm with the centralized version using CPLEX. Examining the SAEV dispatch policy over the first three hours of operation in Figures 6 and 7, we can observe the process of queue formation and energy service. After a simulation window of three hours (solving the dispatch problem every 5 minutes), both policies have stable passenger queues and serve approximately the same portion of the energy demand over time. These plots show that the decentralized policy is able to approximate near-optimal solutions without high levels of communication. Though the solutions each timestep are not always exactly the same (since the ADMM approach is a heuristic), the maximum stability dispatch ensures that these errors are more likely to be corrected at later timesteps [30]. Additional heuristics built into the end of the process, ensuring vehicles are discharging at higher rates to serve all demand could be explored in the future to speed the convergence of the algorithm and further reduce the difference between the two curves in Figure 7.
@@
We run similar tests on the larger Sioux Falls transportation network with the IEEE 85-node distribution network. These networks are shown in Figure 8 and are connected by four sets of charging stations. We test the operations of vehicles when used as SAV, TESSs, and SAEVs and there are two breaks in the electric grid between nodes 31–32 and nodes 34–44. Centralized solutions times range from about 10 minutes to several hours using CPLEX (if they can be found at all). Near-optimal solutions can be found in 1–5 minutes using the decentralized ADMM approach (assuming vehicles can optimize their own routes in parallel). Figure 9 shows the convergence process of the algorithm on the larger Sioux Falls network for the first timestep. This is the most difficult timestep to solve since the few vehicles in the network mean that the greedy solution of each vehicle will lead to many conflicts. However, the primal and dual residuals both drop below $10^{-3}$ in about 60 seconds (about 4 upper-level iterations and 143 lower level iterations). The residuals could be reduced in the future or convergence rates could be increased with additional parameter tuning, though this may be unnecessary given the already fast convergence and the stable, near-optimal solution quality. Additional speed gains can be found in the future by developing heuristics for the lower-level routing problems.
@@
Figure 9: Validation of ADMM algorithm for the Sioux Falls network (points are lower-level iterations and lines are upper-level iterations.
Figures 10 and 11 show a very similar pattern as for the smaller test network. The SAEVs are able to stabilize the demand with only slightly longer queues than if they were operated as SAVs, while also providing power to impacted areas. This is because vehicles already traveling between nodes carrying passengers can stop for a short time to discharge before carrying passengers the other direction and recharging. While neither fleet stabilizes passenger demands in the first 3 hours, the SAEV dispatch shows only slightly longer queues than that SAV fleet.
@@
These case studies suggest several implications for practical applications of SAEV routing methodologies and for future algorithmic improvements. This study suggests that ignoring interactions between SAEVs and the grid can be detrimental as charging and discharging behaviors add additional demands on the time of these vehicles. Ignoring such interactions could lead to underestimates of queue lengths as shown in Figures 4 and 10. On the other hand, ignoring the need to serve passengers (particularly during disaster scenarios) could lead to overestimates of the amount of power SAEVs can carry across outages (see figures 5 and 11). Though this paper focused on disaster scenarios, vehicles may also use vehicle to grid charging to reduce demand fluctuations, provide peak shaving, or many other grid services. However, demand fluctuations of energy and vehicles are linked, so this behavior needs to be better analyzed in a comprehensive framework to understand these impacts.
@@ V Conclusions
Future work should develop heuristics or use learning-based optimizers to solve the problem more quickly (for real-time implementation on realistic networks) and compare the ADMM heuristic to other solution approaches such as Benders decomposition. The lower-level problems of the ADMM heuristic were solved with CPLEX, though many exhibit structures that could be exploited in additional exact or heuristic solution approaches. Additional research should also examine the sensitivity to the cost functions, fleet size, charging/discharging rate, battery size, length of the time horizon, and behavior under different loading scenarios (i.e. more realistic demand profiles). Additional modifications to the optimization problem could include congestion created by the SAEV fleet or investigate the impacts of other energy storage solutions. Similarly, additional objective functions should be tested for day-to-day operation that could minimize demand fluctuations, prioritize green energy sources, incorporate equity concerns, etc. In addition to better understanding the SAEV routing problem, we believe that our proposed ADMM heuristic could be applicable to other vehicle routing approaches. This subproblem was shown to be very effective when the network was congested, and variants could be applied to other large-scale dispatch and routing problems. Finally, we have assumed a cooperative framework between the SAEV dispatcher and the power grid operator, possibly motivated by economic incentives or regulatory frameworks for post-disaster recovery. Future research should explore non-cooperative scenarios, potentially using game-theoretic frameworks to model strategic interactions and design appropriate incentive mechanisms.